OpenStudy (anonymous):

Vectors [2,5,a] and [-6,b,9] are collinear. What is the value of b/a?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

if they are the same vector, divide like parts

OpenStudy (zzr0ck3r):

2/-6 = 5/b

OpenStudy (amistre64):

[2, 5,a] [-6,b,9] the ration is -1/3

OpenStudy (amistre64):

or [-6,b,9] [2, 5,a] the ratio os now -3 :)

OpenStudy (amistre64):

b/5 = 9/a = -3 can you find b and a?